OpenStudy (anonymous):

What is the equivalent resistance for a parallel circuit that has 15V ,R1 =9.1kohms,R2=1.2kohms, R3=6.2kohms

OpenStudy (anonymous):

\(\frac{1}{R_{eq}}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}=1.1 \implies R_{eq}=\frac{1}{1.1}=0.91\)